OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5). Write the equation of the line parallel to the line 4x + 2y = 5 and passing through the point (-3,5). Write the equation of the line parallel to the line 4x + y = -1 and passing through the point (5,0). Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4). Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5). A) y = -5x + 15 B) y = -3x + 14 C) y = -2x - 1 D) y = -4x + 20

OpenStudy (anonymous):

oh, these.....um....i can really only tell u how to do it..im not good at these

OpenStudy (nubeer):

x-3y = 9 change it into general form format which is y =mx +c, so re-arrange your equation x-9= 3y.. now divide both sides by 3.

OpenStudy (anonymous):

i would really like to help, but these i have no idea how to do :(

OpenStudy (nubeer):

y = (x/3) - 3 , now compare it with y = mx + c m= gradient.

OpenStudy (nubeer):

so what would be the gradient of this line @johana ?

OpenStudy (anonymous):

they are matching problems

OpenStudy (nubeer):

ya i figured that up but you still have to find the equation of each line of your question.

OpenStudy (nubeer):

hmm so?

OpenStudy (anonymous):

idk how to do it ??? can u show me step by step

OpenStudy (nubeer):

for 1st x-3y = 9 change it into general form format which is y =mx +c, so re-arrange your equation x-9= 3y.. now divide both sides by 3. y = (x/3) - 3 , now compare it with y = mx + c m= gradient. can you tell me by comparing what is m now?

OpenStudy (compassionate):

For the last one: Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5). Perpendicular lines have opposite slopes. Passing through the points (2, 5) Plug that into Point-Slope form. y - 5 = 1/5(x - 2) y - 5 = 1/5x - 1 y = 1/5x + 4 <-- Equation.

OpenStudy (anonymous):

oh ok i got it how to do it thank you so much :))

OpenStudy (compassionate):

Thanks, nubeer, for helping her, too.

OpenStudy (anonymous):

:))